The pH of solution formed by mixing 40 ml of 0.1M HCl with 10 ml 0.45 M of NaOH is ? Please answer quickly 

Grade:

1 Answers

Sahani Kumar
5 years ago
Concentration of HCl= (0.1*40)/1000= 0.004 mol/L
Concentration of NaOH= (0.45*10)/1000 = 0.0045 mol/L
0.004 mol/L of HCl will produce 0.004 mol/L H+ ions. 
0.0045 mol/L of NaoH will produce 0.0045 mol/L
OH- ions.
But 0.004 H+ ion will netralise0.004 OH-ion.
Thus only 0.0005 mol/L OH- ion will be left free
after neutralisation and only these only will
contribute to pH of the solution. 
Now, -pOH=log[OH-]      -pOH=log[5 * 10^-4]
       -pOH=log5+ log(10^-4)
      -pOH=0.699 – 4.            pOH=4 – 0.699
       pOH=3.301
       pH=pKw – pOH.     pH=14 – 3.301
       pH=10.699